# Random Sperner lemma and random Brouwer fixed point theorem

### v3

@@ Abstract
In this paper, we systematically develop the theory of $L^{0}$-simplicial subdivisions of $L^{0}$-simplexes by grasping the inherent connection between $L^{0}$-simplexes and usual simplexes. We establish a key representation theorem of a proper $L^{0}$-labeling function by usual proper labeling functions. Based on the representation theorem, we can establish a general random Sperner lemma. Next, as an application of the random Sperner lemma, we give a complete proof of the known random Brouwer fixed theorem in random Euclidean spaces. Besides, all the works currently available and closely related to the randomization of the classical Brouwer fixed point theorem are unified.
@@
Random Sperner Lemma, Random Brouwer fixed point theorem, Random Euclidean space, $L^{0}$-module, $L^{0}$-simplex
@@ 1 Introduction
Throughout, $\mathbb{R}^{d}$ denotes the $d$-dimensional Euclidean space. The well known Brouwer fixed point theorem [3] states that every continuous mapping from a bounded closed convex subset of $\mathbb{R}^{d}$ to itself has a fixed point. Brouwer fixed point theorem has rich applications in analysis, geometry, differential equations and mathematical economics [39, 8, 14, 40], in particular, it is also attractive that Brouwer fixed point theorem is equivalent to many famous fundamental results [38, 29, 30, 41], for example, the Knaster-Kuratowski-Mazurkiewicz (briefly, $KKM$) mapping principle [33], the classical Borsuk theorem [2] and the Gustafson-Schmitt-Cid-Mawhin’s theorem on the existence of periodic solutions of ordinary differential equations [8, Theorem 1.1]. Due to its fundamental importance, Brouwer fixed point theorem has been generalized in various ways [32, 46, 50, 7, 13] and has had a major impact on the development of nonlinear analysis and fixed point theory [1, 9, 18].
@@
With the aim of developing random functional analysis [23, 26], dynamic Nash equilibrium theory [11, 26] and stochastic differential equations [42, 43, 44, 45], there are strong desires to extend Brouwer fixed point theorem to the setting of a random Euclidean space $L^{0}(\mathcal{F},\mathbb{R}^{d})$, where $L^{0}(\mathcal{F},\mathbb{R}^{d})$ is the linear space of equivalence classes of random vectors from a probability space $(\Omega,\mathcal{F},P)$ to the $d$-dimensional Euclidean space $\mathbb{R}^{d}$. Since $L^{0}(\mathcal{F},\mathbb{R}):=L^{0}(\mathcal{F},\mathbb{R}^{1})$ is an algebra over the real number field $\mathbb{R}$, $L^{0}(\mathcal{F},\mathbb{R}^{d})$ is a free module of rank $d$ over the algebra $L^{0}(\mathcal{F},\mathbb{R})$, and thus the notions of an $L^{0}$-convex set and a $\sigma$-stable set can be introduced and developed. More importantly, $L^{0}(\mathcal{F},\mathbb{R}^{d})$ can be regarded as a special complete random normed module. The notion of a random normed module was introduced independently by Guo in connection with the idea of randomizing space theory [19, 21, 22, 23, 25, 26] and by Gigli in connection with nonsmooth differential geometry on metric measure spaces [16] (see also [4, 5, 6, 17, 36, 37] for the related advances). When regarded as a random normed module, $L^{0}(\mathcal{F},\mathbb{R}^{d})$ is often endowed with two kinds of useful topologies: one is the $(\varepsilon,\lambda)$-topology, denoted by $\mathcal{T}_{\varepsilon,\lambda}$, which coincides with the usual topology of convergence in probability, is a typical locally nonconvex linear topology and makes $(L^{0}(\mathcal{F},\mathbb{R}^{d}),\mathcal{T}_{\varepsilon,\lambda})$ a topological module over the topological algebra $(L^{0}(\mathcal{F},\mathbb{R}),\mathcal{T}_{\varepsilon,\lambda})$; the other is the locally $L^{0}$-convex topology, denoted by $\mathcal{T}_{c}$, which was introduced in [15] and coincides with the topology generated by random open balls and makes $(L^{0}(\mathcal{F},\mathbb{R}^{d}),\mathcal{T}_{c})$ a topological module over the topological ring $(L^{0}(\mathcal{F},\mathbb{R}),\mathcal{T}_{c})$, see [23] for the relevant connections between the two kinds of topologies. Thus, according to different choices of $\mathcal{T}_{\varepsilon,\lambda}$ and $\mathcal{T}_{c}$, random Brouwer fixed point theorem has the corresponding two versions (or forms): the first version is the formulation of random Brouwer fixed point theorem under $\mathcal{T}_{c}$, which was first given in [26] and is stated as Theorem 2.1 in this paper; the second version is the formulation of random Brouwer fixed point theorem under $\mathcal{T}_{\varepsilon,\lambda}$, which was first given in [11, 28] and is stated as Proposition 2.3 in this paper. It follows immediately from Lemma 4.3 of [26] that Proposition 2.3, namely the $\mathcal{T}_{\varepsilon,\lambda}$-version of random Brouwer fixed point theorem is a special case of Theorem 2.1 (namely the $\mathcal{T}_{c}$-version of random Brouwer fixed point theorem). Hence, from now on, Theorem 2.1 is called random Brouwer fixed point theorem.
The focus of [26] lies in the proof of a noncompact Schauder fixed point theorem in random normed modules, where the authors merely presented a sketch proof of Theorem 2.1 based on the works [11, 28]. Considering the fundamental importance of Theorem 2.1, one of the purposes of this paper is to give a complete proof of Theorem 2.1 in order to make the work [26] self-contained. It is in the process of achieving the proof that many basic problems are treated thoroughly.
It is well known that the proof of the classical Brouwer fixed point theorem is not very easy and the simplest of all the classical proofs is by means of Sperner lemma [1, 33, 35, 48]. The classical Sperner lemma states that for any simplicial subdivision $\mathscr{S}$ of a simplex and any proper labeling function of $\mathscr{S}$ there are an odd number of completely labeled simplexes in $\mathscr{S}$. It is also well known that the most accessible proof of Sperner lemma was given by Kuhn in [35] via the graph-theoretic method. However, the proof of random Brouwer fixed point theorem involves more since $L^{0}(\mathcal{F},\mathbb{R}^{d})$ admits a more complicated topological and algebraic structure than $\mathbb{R}^{d}$, which requires the following three nontrivial tools. The first tool is Lemma 4.3 of [26], which states that a $\sigma$-stable mapping is $\mathcal{T}_{c}$-continuous iff it is random sequentially continuous. Since the topology $\mathcal{T}_{c}$ is too strong for us to conveniently treat the fixed point problem of the $\mathcal{T}_{c}$-continuous mappings, but random sequential continuity can be described in terms of $\mathcal{T}_{\varepsilon,\lambda}$, this lemma enables us to make full use of $\mathcal{T}_{\varepsilon,\lambda}$ to treat random Brouwer fixed point theorem. The second tool is the randomized version of the classical Bolzano-Weierstrass theorem [31], which states that an almost surely (briefly, a.s.) bounded sequence in $L^{0}(\mathcal{F},\mathbb{R}^{d})$ always admits an a.s. convergent randomized subsequence although it may admit no a.s. convergent subsequence. Since an a.s. bounded, $\mathcal{T}_{\varepsilon,\lambda}$-closed and $L^{0}$-convex subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$, is generally noncompact but random sequentially compact by the randomized Bolzano-Weierstrass theorem, then the randomized Bolzano-Weierstrass theorem can be used to overcome the difficulty due to noncompactness [11, 26, 28]. The third tool is a random version of the classical Sperner lemma. To prove a special case of the $\mathcal{T}_{\varepsilon,\lambda}$-version of random Brouwer fixed point theorem (see [11, Proposition 3.1]), the authors [11] first give a special case of random Sperner lemma (see the first part of [11, Theorem 2.3]), where the authors studied a special $L^{0}$-simplicial subdivision of an $L^{0}$-simplex (called the $L^{0}$-barycentric subdivision) in detail. The special random Sperner lemma provided in [11] was employed in all the works [11, 26, 28] on random Brouwer fixed point theorem. However, the proof of [11] on the special random Sperner lemma is not easily accessible to the reader, for example, the idea of the proof in [11] was very similar to Kuhn’s one [35], but the method in [11] was not by means of the graph-theoretic argumentation and thus rather complicated. Considering the fundamental importance of random Brouwer fixed point theorem, for example, it has been used in [26] for the proof of a noncompact Schauder fixed point theorem in random normed modules, and it will also be used to establish a noncompact $KKM$ principle in random locally convex modules so that a series of open problems posed in [26] will be solved in our forthcoming work, it is necessary to look for a very easily accessible proof of random Sperner lemma and to further give a complete proof of random Brouwer fixed point theorem. This leads us directly to systematically developing the theory of $L^{0}$-simplicial subdivisions of $L^{0}$-simplexes.
In fact, our approach to developing $L^{0}$-simplicial subdivisions is completely new, in particular our idea of defining an $L^{0}$-simplicial subdivision of an $L^{0}$-simplex is directly linking the $L^{0}$-simplicial subdivision to the corresponding counterpart simplicial subdivision of the corresponding counterpart simplex, see Definition 2.6. An advantage of Definition 2.6 enables us to establish a key representation theorem of a proper $L^{0}$-labeling function by usual proper labeling functions, see Theorem 3.5 of this paper for details. Consequently, by using the representation theorem we can smoothly establish a more general random Sperner lemma than the one given in [11], namely Theorem 2.9, in particular its proof is very easily accessible to the reader since this proof will be eventually finished by directly applying the classical Sperner lemma. Finally, combining Theorem 2.9 and the idea of the proof of Lemma 4.6 of [26], we can further give a complete proof of random Brouwer fixed point theorem.
@@
Recently, we have known from [43, 44, 45] that Ponosov has been working on a fixed point theory for a class of local and continuous-in-probability operators with the aim of providing a tool for the study of stochastic differential equations. In particular, in [44] Ponosov proved a stochastic version of Brouwer fixed point theorem, which is stated as Proposition 2.4 in this paper and has played an essential role in the proofs of the two fixed point theorems — Theorems 5.1 and 5.2 of [45] that can be directly applied to the study of stochastic differential equations. In this paper we prove that Ponosov’s stochastic Brouwer fixed point theorem [44] is equivalent to the $\mathcal{T}_{\varepsilon,\lambda}$-version of random Brouwer fixed point theorem by combining Ponosov’s deep result [42] and the theory of complete random inner product modules. In particular, we further obtain a somewhat surprising conclusion again by means of Ponosov’s result [42]: a $\mathcal{T}_{\varepsilon,\lambda}$-continuous $\sigma$-stable mapping from a $\mathcal{T}_{\varepsilon,\lambda}$-closed $L^{0}$-convex subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ to $L^{0}(\mathcal{F},\mathbb{R}^{d})$ must be a.s. sequentially continuous (see Proposition 5.8 of this paper), which shows that Proposition 3.1 of [11] is equivalent to either of Proposition 2.3 and Proposition 2.4.
@@
Since the classical Borsuk theorem [2] is of fundamental importance, which states that the unit sphere of an Euclidean space is not a retraction of the closed unit ball, and is also equivalent to classical Brouwer fixed point theorem. In [11], the $\mathcal{T}_{\varepsilon,\lambda}$-version of random Borsuk theorem was first proved, see Lemma 3.5 of [11], but the equivalence of the $\mathcal{T}_{\varepsilon,\lambda}$-version of random Borsuk theorem with the $\mathcal{T}_{\varepsilon,\lambda}$-version of random Brouwer fixed point theorem had been not yet given. In the end of this paper we not only prove the equivalence of the $\mathcal{T}_{\varepsilon,\lambda}$-version of random Borsuk theorem with the $\mathcal{T}_{\varepsilon,\lambda}$-version of random Brouwer fixed point theorem but also prove a general random Borsuk theorem and together with its equivalence with random Brouwer fixed point theorem.
@@
This paper is organized as follows. In Section 2, we provide some prerequisites and a survey of different forms and special cases of random Brouwer fixed point theorem currently available. In Section 3, we establish a key representation theorem of a proper $L^{0}$-labeling function by usual proper labeling functions, namely Theorem 3.5. Based on Theorem 3.5, we prove Theorem 2.9, namely the random Sperner lemma. Next, as an application of Theorem 2.9, in Section 4, we prove Theorem 2.1. In Section 5, we prove the equivalence of Proposition 2.3 and Proposition 2.4. By the way, we also give some other interesting results in Section 5, which are closely related to the topic on random Brouwer fixed point theorem. Finally, in Section 6, we prove random Borsuk theorem — Theorem 6.4 is equivalent to random Brouwer fixed point theorem.
@@
2 Random Brouwer fixed point theorem and Random Sperner lemm. Prerequisites
2.1 Different forms and special cases of random Brouwer fixed point theorem
To clearly present the main contents of this section, we first introduce some basic notation and terminologies.

### v5

@@ Abstract
We first establish a general random Sperner lemma by presenting a completely new approach for the theory of $L^{0}$-simplicial subdivisions of $L^{0}$-simplexes. Based on this, we are able to achieve a new complete proof of the random Brouwer fixed theorem in random Euclidean spaces, which can provide a solid foundation for various contemporary applications of interest. Afterward, we unify the works currently available and closely related to the random Brouwer fixed theorem: we first prove that the stochastic Brouwer fixed point theorem occurring elsewhere in stochastic analysis is equivalent to a special case of our random Brouwer fixed theorem, and then prove a general random Borsuk theorem and its equivalence with the random Brouwer fixed theorem. Finally, we conclude this paper with commentaries on recent state of study of the famous Schauder conjecture.
@@
Random Euclidean space, $L^{0}$-simplex, Random Sperner Lemma, Random Brouwer fixed point theorem, Random Borsuk theorem
@@ 1 Introduction
The celebrated Brouwer fixed point theorem [3] states that every continuous map from a bounded closed convex subset of $\mathbb{R}^{d}$ to itself has a fixed point. Here and throughout, $\mathbb{R}^{d}$ denotes the $d$-dimensional Euclidean space. The Brouwer fixed point theorem has numerous applications in analysis, geometry, differential equations and mathematical economics [41, 9, 15, 42]. Furthermore, it is also attractive that Brouwer fixed point theorem is equivalent to many prominent fundamental results [40, 30, 31, 43], for example, the Knaster-Kuratowski-Mazurkiewicz (briefly, $KKM$) mapping principle [34], the classical Borsuk theorem [2] and the Gustafson-Schmitt-Cid-Mawhin theorem on the existence of periodic solutions of ordinary differential equations [9, Theorem 1.1]. Due to its fundamental importance [7], the Brouwer fixed point theorem has been given many proofs [41, 43]. One of the most elegant proofs of all of its classical proofs is via the Sperner lemma [51], in particular that proof also helps in developing a simplicial approximation of fixed points [36, 48]. Further, the Brouwer fixed point theorem has been generalized in various ways [33, 49, 54] and has had a major impact on the development of nonlinear analysis and the general fixed point theory [1, 10, 19, 8, 14, 56].
@@
With an aim of developing random functional analysis [24, 27], dynamic Nash equilibrium theory [12, 27] and stochastic differential equations [44, 45, 46, 47], there have been strong desires to extend the Brouwer fixed point theorem to the setting of a random Euclidean space $L^{0}(\mathcal{F},\mathbb{R}^{d})$, where $L^{0}(\mathcal{F},\mathbb{R}^{d})$ is the linear space of equivalence classes of random vectors from a probability space $(\Omega,\mathcal{F},P)$ to the $d$-dimensional Euclidean space $\mathbb{R}^{d}$. Since $L^{0}(\mathcal{F},\mathbb{R}):=L^{0}(\mathcal{F},\mathbb{R}^{1})$ is an algebra over the real number field $\mathbb{R}$, $L^{0}(\mathcal{F},\mathbb{R}^{d})$ is a free module of rank $d$ over the algebra $L^{0}(\mathcal{F},\mathbb{R})$, and thus the notions of an $L^{0}$-convex set and a $\sigma$-stable set can be introduced and developed. More importantly, $L^{0}(\mathcal{F},\mathbb{R}^{d})$ can be regarded as a special complete random normed module. The notion of a random normed module was introduced independently by Guo [20, 22, 23, 24, 26, 27] in connection with the idea of random-structured space theory and by Gigli [17] in connection with nonsmooth differential geometry on metric measure spaces (see also [4, 6, 5, 18, 37, 38] for related advances). When regarded as a random normed module, $L^{0}(\mathcal{F},\mathbb{R}^{d})$ is often endowed with two topologies: one is the $(\varepsilon,\lambda)$-topology, denoted by $\mathcal{T}_{\varepsilon,\lambda}$, which coincides with the usual topology of convergence in probability and is a typical metrizable locally nonconvex linear topology and makes $(L^{0}(\mathcal{F},\mathbb{R}^{d}),\mathcal{T}_{\varepsilon,\lambda})$ a topological module over the topological algebra $(L^{0}(\mathcal{F},\mathbb{R}),\mathcal{T}_{\varepsilon,\lambda})$, while the other is the locally $L^{0}$-convex topology, denoted by $\mathcal{T}_{c}$, which was introduced in [16], coincides with the topology generated by random open balls, and makes $(L^{0}(\mathcal{F},\mathbb{R}^{d}),\mathcal{T}_{c})$ a topological module over the topological ring $(L^{0}(\mathcal{F},\mathbb{R}),\mathcal{T}_{c})$; see [24] for the inherent connections between the two kinds of topologies. Thus, according to different choices of $\mathcal{T}_{\varepsilon,\lambda}$ and $\mathcal{T}_{c}$, a random Brouwer fixed point theorem has possibly two corresponding formulations: the first is the formulation of random Brouwer fixed point theorem equipped with $\mathcal{T}_{c}$, which was first given in [27] and is stated as Theorem 2.1 in this paper, while the second is the formulation with $\mathcal{T}_{\varepsilon,\lambda}$, which was first given in [12, 29] and is stated as Proposition 2.3 in this paper. It follows immediately from Lemma 4.3 of [27] that Proposition 2.3, namely, the $\mathcal{T}_{\varepsilon,\lambda}$-version of a random Brouwer fixed point theorem is a special case of Theorem 2.1 (namely, the $\mathcal{T}_{c}$-version of random Brouwer fixed point theorem). Hence, from now on, Theorem 2.1 will be called by us the random Brouwer fixed point theorem.
Just as the Brouwer fixed point theorem is the cornerstone of the topological fixed point theory, the random Brouwer fixed point theorem is of fundamental importance, for example, it has been used in [27] for the proofs of both a noncompact Schauder fixed point theorem in random normed modules and the existence theorem of a conditional Nash equilibrium point. In particular, it will also be used to solve the open problems posed in [27], for example, in a forthcoming work where we will establish a noncompact $KKM$ mapping principle and further prove a general dynamic Nash equilibrium theorem while we will also provide an effective approximation algorithm for searching the dynamic Nash equilibrium point. The original proof of the random Brouwer fixed point theorem [27] essentially employed a special random Sperner lemma (see the first part of [12, Theorem 2.3]). However, the special random Sperner lemma has obvious limitations in its applications, and its proof given in [12] is also inaccessible to the reader, which renders our original proof of the random Brouwer fixed point theorem in [27] neither complete nor satisfactory. Therefore, one of the central purposes of this paper is to give a new complete proof of the random Brouwer fixed point theorem. For this sake, in the frontal part of this paper we will first provide a new foundation (namely, a general random Sperner lemma) for such a complete proof.
To emphasize the necessity of reconsidering the problem of randomizing the classical Sperner lemma, let us first recall the classical Sperner lemma and its remarkable applications, and give a brief comment on the previous work [12] on a special random Sperner lemma. In 1928, this classical lemma was given by Sperner [51], also stated as Lemma 2.8 in this paper for the reader’s convenience. Four decades later in 1968 an elegant proof was given by Kuhn [36] via a graph-theoretic argument. As has been mentioned above, the Sperner lemma-based proof of the classical Brouwer fixed point theorem is significant since it has the following two merits. One is its elegance, the other is the constructiveness that leads to a simplicial approximation of fixed points [48, 36]. Mainly motivated by the proof of the classical Brouwer fixed point theorem, Drapeau et al. [12] first gave a special random Sperner lemma (see the first part of [12, Theorem 2.3]) and further proved a special case of the random Brouwer fixed point theorem on an $L^{0}$-simplex (see the second part of [12, Theorem 2.3]). Just as stated above, the classical Sperner lemma is so fundamental and important that it merits a complete random version, but the related, existing work in [12] has certain significant limitations, listed as follows:
  1. (1)
    Concerning the $L^{0}$-simplicial subdivision of an $L^{0}$-simplex, limited by their method authors of [12] only considered the case of the $L^{0}$-barycentric subdivision so that the special random Sperner lemma did not encompass the classical Sperner lemma as a special case.
  2. (2)
    The $L^{0}$-barycentric subdivision is too special to ensure that the higher-fold $L^{0}$-barycentric subdivision of an $L^{0}$-simplex is again an $L^{0}$-barycentric subdivision so that their notion of an $L^{0}$-labeling function [12, Definition 2.1] is not as natural as that of the usual labeling function. When the special random Sperner lemma was applied to the proof of their special random Brouwer fixed point theorem on an $L^{0}$-simplex [12, Theorem 2.3] the so obtained sequence of completely labeled $L^{0}$-simplexes approaching the desired fixed point is not decreasing, which rules out the possibility of designing an effective algorithm for the simplicial approximation of fixed points.
  3. (3)
    Their proof of the special random Sperner lemma was similar to that of Kuhn [36] for the classical Sperner lemma, but authors of [12] did not adopt a graph-theoretic argument and, thus, their proof was rather complicated, lacking readability as well as verifiability.
To establish a more appeasing random Sperner lemma in order to overcome the limitations of the earlier approach , we first present a completely new approach for the theory of $L^{0}$-simplicial subdivisions of $L^{0}$-simplexes. In Sections 2, 3 and 4, we introduce a general notion of an $L^{0}$-simplicial subdivision of an $L^{0}$-simplex in connection with a usual simplicial subdivision of a classical simplex (see Definition 2.6), which helps us establish a key representation theorem of a proper $L^{0}$-labeling function by usual proper labeling functions (see Theorem 3.5). In turn, this representation theorem can help build a general random Sperner lemma in a concise way as a consequence of the classical Sperner lemma, where our success is achieved through the connection between proper $L^{0}$-labeling functions and usual proper labeling functions rather than by considering the classification of $L^{0}$-subsimplexes as in the proof of the special random Sperner lemma [12]. The general random Sperner lemma not only includes the classical Sperner lemma as a special case but also meets all the needs of our future applications. Further, by combining the general random Sperner lemma, Lemma 4.3 of [27] and the randomized Bolzano-Weierstrass theorem [32] we are able to give a new complete proof of the random Brouwer fixed theorem in random Euclidean spaces and, in particular, we note that the proof contains constructive information that is quite desirable (see Remark 4.11).
@@
Recently, we have known from [45, 46, 47] that Ponosov has been working on a fixed point theory for a class of local and continuous-in-probability operators with the aim of providing a tool for the study of stochastic differential equations. In particular, in [46] Ponosov proved a stochastic version of Brouwer fixed point theorem, which is stated as Proposition 2.4 in this paper and has played an essential role in the proofs of the two fixed point theorems — Theorems 5.1 and 5.2 of [47] that can be directly applied to the study of stochastic differential equations. In Section 5, we begin in proving that Ponosov’s stochastic Brouwer fixed point theorem [46] is equivalent to a special case of our random Brouwer fixed point theorem by combining Ponosov’s profound result [44] and the theory of complete random inner product modules. In particular, we are able to obtain a somewhat surprising conclusion again by means of Ponosov’s result [44]: a $\mathcal{T}_{\varepsilon,\lambda}$-continuous $\sigma$-stable mapping from a $\mathcal{T}_{\varepsilon,\lambda}$-closed $L^{0}$-convex subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ to $L^{0}(\mathcal{F},\mathbb{R}^{d})$ must be a.s. sequentially continuous (see Proposition 5.8 of this paper), which shows that Proposition 3.1 of [12] is equivalent to either Proposition 2.3 or Proposition 2.4.
@@
The classical Borsuk theorem [2] states that the unit sphere of an Euclidean space is not a retraction of the closed unit ball, and in turn is also equivalent to the classical Brouwer fixed point theorem. In [12], a special random Borsuk theorem was proved, see Lemma 3.5 of [12]. In this paper we are able to prove a general random Borsuk theorem, and in particular we also prove that it is also equivalent to the random Brouwer fixed point theorem.
@@
This paper is organized as follows. In Section 2, we provide some prerequisites and a survey of the random Brouwer fixed point theorem and its special cases currently in existence. In Section 3, we establish a key representation theorem of a proper $L^{0}$-labeling function by usual proper labeling functions, namely, Theorem 3.5. Based on Theorem 3.5, we prove Theorem 2.9, namely, a random Sperner lemma. As an application of Theorem 2.9, in Section 4, we give a new complete proof of Theorem 2.1. In Section 5, we prove the equivalence of Proposition 2.3 and Proposition 2.4. In Section 6, we prove a general random Borsuk theorem — Theorem 6.4 together with its equivalence with the random Brouwer fixed point theorem. Finally, in Section 7 we conclude this paper with commentaries on recent state of study of the famous Schauder conjecture.
@@
2 The random Brouwer fixed point theorem and a general random Sperner lemm. Preliminaries
2.1 The random Brouwer fixed point theorem and its special cases
We begin by first introducing some basic notation and terminologies.